# HarmonRank: Ranking-aligned Multi-objective Ensemble for Live-streaming E-commerce Recommendation

### v2

@@ Abstract.
Recommendation for live-streaming e-commerce is gaining increasing attention due to the explosive growth of the live streaming economy. Different from traditional e-commerce, live-streaming e-commerce shifts the focus from products to streamers. This shift requires the ranking mechanism to balance both purchases (short-term value) and user-streamer interactions (long-term value), e.g., follows and comments, for the ecology health of the platform. To trade off multiple objectives, a popular solution is to build an ensemble model to integrate multi-objective scores into a unified score. The ensemble model is usually supervised by multiple independent binary classification losses of all objectives. However, this paradigm suffers from two inherent limitations. First, the optimization direction of the binary classification task is misaligned with the ranking task (evaluated by AUC). Second, this paradigm overlooks the alignment between objectives, e.g., comment and buy behaviors are partially dependent which can be revealed in labels correlations. The model can achieve better trade-offs if it learns the aligned parts of ranking abilities among different objectives.
To mitigate these limitations, we propose a novel multi-objective ensemble framework HarmonRank to fulfill both alignment to the ranking task and alignment among objectives. For alignment to ranking, we formulate ranking metric AUC as a rank-sum problem and utilize differentiable ranking techniques for ranking-oriented optimization. For inter-objective alignment, we change the original one-step ensemble paradigm to a two-step one. In 1st step, we align each objective with other objectives in a self-attention mechanism. In 2nd step, we fuse aligned objective encodings into the ensemble score with the guidance of personalized features.
Extensive offline experiments results on two industrial datasets and online experiments demonstrate that our approach significantly outperforms existing state-of-the-art methods. The proposed method has been fully deployed in Kuaishou’s live-streaming e-commerce recommendation platform with 400 million DAUs, contributing over 2% purchase gain.
@@ 1. Introduction
Figure 1. Conceptual view of the multi-objective ensemble stage and different ensemble paradigms. (a) The position of multi-objective ensemble stage in industrial recommender systems. (b) Existing one-step ensemble paradigm based on multi-layer non-linear with multiple binary classification losses. (c) Limitations of existing paradigm. First, lack of the alignment to ranking task. Second, lack of the alignment between objectives.(d) Our proposed paradigm to fulfill alignment to ranking and inter-objective alignments.
Live-streaming e-commerce market has recently experienced explosive growth. Different from traditional e-commerce, the products are endorsed and promoted by streamers in live-streaming e-commerce, which shifts the focus from products to the streamer. This paradigm shift necessitates a change from purchase-oriented ranking mechanism to multi-objective balanced mechanism including interactions between users and streamers, e.g., follows and comments. Different from short-term purchase, these interaction-oriented objectives matter to the ecosystem for the their long-term value. For an example, follow behaviors can foster long-term user-streamer relationships, which helps to the ecology health of the platform. This change poses more challenges for balancing these competing objectives than traditional purchase-driven e-commerce recommender systems.
To balance these competing objectives, industrial ranking systems usually follow a two-stage paradigm (see Fig. 1 (a)), i.e., multi-objective prediction (MP) and multi-objective ensemble (ME) (Pei et al., 2019; Cao et al., 2025; Li et al., 2023; Carmel et al., 2020). The MP stage is usually a high-capacity multi-task learning model (Ma et al., 2018a) whose goal is to predict the precise probability scores for multiple objectives, e.g., click through rate (CTR). Whereas ME stage is a light-weight model whose goal is to generate a unified ensemble score in a personalized manner. It usually only takes aforementioned multi-objective scores and a few user features as inputs. The ensemble score is used to truncate top items to represent for users. In this paper, we focus on ME problem. Similar to MP stage, in ME stage we use AUC (Area Under ROC Curve) (Bamber, 1975) between the ensemble score and the ground-truth labels to evaluate the ranking abilities of the ensemble model. The only difference is that we utilize the sum of AUC of all objectives as the metic for comprehensive multi-objective ranking evaluation.
The key challenge at this stage lies in the absence of a gold-standard supervision to comprehensively assess items considering multiple objectives. To mitigate this problem, a typical solution is to impose multiple independent binary classification supervisions over the ensemble score (Cao et al., 2025; Lin et al., 2019) (see Fig. 1 (b)). However, this paradigm suffers from two limitations, for which we present a conceptual view in Fig. 1 (c). (1) Lack of alignment to ranking task. The optimization directions of binary classification based supervisions are misaligned with the ultimate ranking task (evaluated by AUC) (Tang et al., 2022). For an example, if the positive labels of some action are highly sparse, a dummy predictor which predicts all instances as negative seems to present an acceptable binary classification result. But it actually yields a poor ranking results with an AUC of 0.5. (2) Lack of inter-objective alignment. The different objectives are neither independent nor strictly adversarial - they are partially dependent. For an example, buy actions empirically exhibit high correlations with comment actions, which can be verified through the statistical analysis of correlation between objectives (see Fig.4). Thus, if model can learn the aligned parts among multi-objectives, it is easier for model to trade-off these competing objectives.
@@
To mitigate these issues, in this paper, we propose a novel personalized multi-objective ensemble framework HarmonRank, which enables both alignment to the ranking task and alignment between multiple objectives. For alignment to ranking, we firstly reformulate the non-differentiable AUC as a rank sum problem. Then we utilize an advanced differentiable ranking technique to optimize multi-objective AUC in an end-to-end manner. For alignment between objectives, we propose to change existing one-step ensemble architecture to a two-step one, i.e., first align then ensemble. In align step, we align the common parts between multiple objectives via self-attention mechanism. In ensemble step, we fuse the aligned encodings of all objectives into ensemble score. In this process, we compress personalized information into a query and enable user-specific ensemble by cross attention mechanism.
@@
  • We propose a novel personalized multi-objective ensemble framework HarmonRank to enable the alignment of optimization to the ranking task with differentiable AUC optimization.
  • We propose a two-step ensemble paradigm to enable align the shared parts between multiple objectives, to obtain better trade-off between objectives.
  • We conduct extensive offline experiments on two industrial recommendation datasets. Experimental results demonstrate that our HarmonRank achieves superior performance over all existing state-of-the-art methods. Additionally, HarmonRank contributes more than 2% purchase gain for Kuaishou live-streaming e-commerce platform.
@@ 4. HarmonRank
4.1. Alignment to ranking
@@ 4.1.2. Differentiable Ranking
4.2. Alignment among objectives
We find prior wisdom consistently overlook the alignment between multiple objectives, which makes model hard to learn the shared ranking abilities among objectives. As shown in Fig. 4, different objectives can be partially aligned with varying degrees. To fill this gap, we design a dual-path architecture to integrate multiple scores in both one-step (relation-agnostic) manner and two-step (relation-aware) manner between objectives. As shown in Fig. 2 (a), we split our framework into two parallel components: relation-aware module and relation-agnostic module. The output scores of two modules are lately fused into a united score to compute the AUC loss.
@@ 4.2.1. Relation-aware Module
We align the shared parts between objectives with a relation-aware two-step paradigm, first align then ensemble. For align step, as shown in Fig. 2(b), we utilize self-attention mechanism (Vaswani et al., 2017) to capture pairwise relations between objectives. Through dynamic computation of relations between objectives, the aligned common parts between them are extracted. Specifically, the softmax normalized dot products between projected objective vectors, $\mathbf{Q}_{r}$ and $\mathbf{K}_{r}$, form the self attention matrix $\mathbf{A}_{r}\in\mathbb{R}^{m\times m}$ between them. This matrix naturally represents the relation strengths between objectives. Then the original score embeddings are integrated into the representations $\mathbf{x}_{r}$ according to the attention matrix.
@@
Although we achieve inter-objective alignment, we still lack clear guidance on how to fuse multiple objectives into a unified score. Fortunately, user-specific personalized information, including user profile and context feature, can serve as strong inductive bias to guide this fusion. We cast the personalized information as a query, and search with it over multi-objective scores. We use the searched result as the fused score encodings to obtain ensemble score. In this way, we implicitly let model answer a question that ‘Which objects can best represent the user’s holistic intentions in the current context?’. Specifically, we transform personalized information into the query vector $\mathbf{Q}_{p}$ by linear projection, similar to key and value vectors.
@@ 5. Experiments
  • RQ1: How does our model perform on industrial recommenda- tion datasets?
  • RQ2: How our module design and hyperparameter choice impact the performance?
  • RQ3: Can our method bring improvements of A/B test metrics on online product environment?
  • RQ4: How our method impact trade-off between objectives? What do the model learn in multiple objective alignment?
@@ 5.1. Implementation Details
Before training ensemble model, we firstly prepare the input multi-objective scores via performing inference with a high-capacity ranking model over both datasets. For Kuaishou-Elive, we use the online in-service model trained incrementally by years of logs. For the public TenRec dataset, we train a standard MMOE (Ma et al., 2018a) ranking model. Based on prepared multi-objective scores, all offline models are trained with 500 epochs with SGD optimizer. In terms of online training model with streaming data, it is usually trained for one epoch for the consideration of run-time cost and one-epoch overfiting risk in ranking stage. This scheme results in low sample efficiency. Since the model of ME stage is much more lightweight than MP stage and no sparse ID feature is used (cause of one-epoch overfitting), we train 20 epochs for each batch of streaming samples, to balance run-time cost and sample efficiency. We set learning rate to 1e-4 and batch size = 10240 for all methods.
@@ 5.2. Datasets
Since there are few public live-streaming e-commerce dataset, we build a new dataset called Kuaishou E-Live, which contains both e-commerce feedback (purchase) and engagement oriented feedbacks (long view, like, follow and comment). We built it by sampling millions of exposure logs at Kuaishou e-commerce platform.
Moreover, to verify the universality of over method, we conduct experiments on a public short video recommendation dataset, TecRec-QKVideo (Yuan et al., 2022). TecRec-QKVideo is a collected from QQ-KAN video feeds platforms. It contains over 1M user interaction logs including five type of feedbacks, click, like, share, follow and long view. For the need of personalized ensemble, we utilize several user profile features like age, gender, etc. We train our method and compared methods with the same equal loss weights (all weights set to 1.0) for fair comparison on objective trade-offs.
@@ 5.5.2. Model components
5.6. Online Results (RQ3)
Figure 4. Consistency analysis between the learned attention matrix $\mathbf{A}_{r}$ and ground-truth data distribution (label’s Spearman Rank Correlation $\rho$.
@@ 5.6. Online Results (RQ3)
Objectives Gain
Purchase +2.635%
Comment +3.034%
Follow +0.451%
Watch time +0.290%
Like +1.673%
@@
5.7. Analysis (RQ4)
@@
5.7.2. Inter-objective alignment analysis.
To shed light on the factual degree of alignment between objectives (ground-truth), and which of them can be learned in inter-objective alignment modules (predicted), we visualize inter-objective correlations in two ways. For the ground-truth relations, we compute Spearman’s rank correlation coefficients $\mathbf{\rho}$ between different objectives over the labels, which measures the monotonic alignment between the ranks of two objective labels. For predicted relations, we visualize the attention weights inferred by our method’s relation module.
(22) $\mathbf{\rho}_{mn}$ $=\text{Spearmanr}(\mathbf{y}_{m},\mathbf{y}_{n})$
(23) $\mathbf{A}_{r}^{{mn}}$ $=\mathbf{A}_{r}[m,n]$
As shown in the Fig. 5, we present the $\rho_{buy,:}$ and $\mathbf{A}_{r}^{buy,:}$ between buy objective and all other objectives. It can be observed that the learned attention matrix exhibits strong alignment with label ranking, showing consistent patterns: comment ¿ long view ¿ follow ¿ like. This demonstrates that our inter-objective alignment module effectively captures latent dependencies. The particularly strong correlations of buy v.s. long view and buy v.s. comment are because the users with immediate purchase willingness tend to engage in prolonged viewing for product explanations and active questioning via comments.
Moreover, to quantitatively validate the consistency between learned patterns in the attention matrix and ground-truth inter-objective alignment, we try to compute the correlation strength between the ground-truth and predicted relations. To preserve more absolute correlation scales in attention matrix, we apply sigmoid instead of softmax normalization over the original dot product matrix. We compute the correlation between label Spearman’s $\mathbf{\rho}$ and predicted attention matrix $\mathbf{A}_{r}$.
(24) $r$ $=\text{Pearson}(\mathbf{\rho},\mathbf{A}_{r})$
Results in Fig .4 demonstrate remarkable consistency ($r=0.82$, $p=0.004$) between learned attentions and empirical alignment, confirming our method’s capability to accurately capture inter-objective alignment by our two-step ensemble paradigm.
Figure 5. Two views of the correlations between buy objective and other objectives: Spearman’s rank correlation ($\mathbf{\rho}$) of labels and learned attention matrix $\mathbf{A}_{r}$.

### v3

# Rethinking Multi-objective Ranking Ensemble in Recommender System: From Score Fusion to Rank Consistency

@@ Abstract.
The industrial recommender systems always pursue more than one business goals, e.g., short term purchases and long term user activeness in e-commerce recommender systems. The inherent intensions between objectives pose significant challenges for ranking stage. A popular solution is to build a multi-objective ensemble (ME) model to integrate multi-objective predictions into a unified score. Although there have been some exploratory efforts, few work has yet been able to systematically delineate the core requirements of ME problem. We rethink ME problem from two perspectives. From the perspective of each individual objective, to achieve its maximum value the scores should be as consistent as possible with the ranks of its labels. From the perspective of the entire set of objectives, an overall optimum can be achieved only when the scores align with the commonality shared by the majority of objectives. However, none of existing methods can meet these two requirements. First, popular methods are optimized by classification or regression tasks, which is misaligned with ranking task. Second, although these methods build complex user-specific structure for multi-objective score fusion, none of them consider modeling the underlying commonality across objectives.
To fill this gap, we propose a novel multi-objective ensemble framework HarmonRank to fulfill both requirements. For rank consistency, we formulate rank consistency (AUC) metric as a rank-sum problem and make the model optimized towards rank consistency in an end-to-end differentiable manner. For commonality modeling, we change the original relation-agnostic ensemble paradigm to a relation-aware one in two steps. First, we capture the commonality among objectives with via self-attention mechanism. Second, we fuse relation-aware objective representations into the ensemble score in a user-specific manner.
Extensive offline experiments results on two industrial datasets and online experiments demonstrate that our approach significantly outperforms existing state-of-the-art methods. Besides, our method exhibits superior robustness to label skew situations which is common in industrial scenarios. The proposed method has been fully deployed in Kuaishou’s live-streaming e-commerce recommendation platform with 400 million DAUs, contributing 2.6% purchase gain.
@@ 1. Introduction
Figure 1. Conceptual view of ranking stage and requirements of an ideal ensemble score. (a) Two sub-stages of ranking stage in industrial recommender systems, i.e., multi-objective prediction and multi-objective ensemble stages. (b) Two requirements of an ideal ensemble score. For each individual objective, the score should align with ranks of labels. For all objectives, the score need align with the majority of them.
Background. The industrial recommender systems always pursue more than one business goals. For instance, e-commerce recommender systems need to balance purchases (short-term gains) and user interactions (long-term ecosystem health). The inherent conflicts among these business goals pose significant challenges for multi-objective ranking stage in these recommender systems, e.g., straightforward pursuing purchase may reduce following behaviors of users, similar to exploit-exploration seesaw problem. To balance these competing objectives, industrial ranking systems usually follow a two-stage paradigm (see Fig. 1 (a)), i.e., multi-objective prediction (MP) and multi-objective ensemble (ME) (Pei et al., 2019; Cao et al., 2025; Li et al., 2023; Carmel et al., 2020). The MP stage is usually a high-capacity multi-task learning (MTL) model (Ma et al., 2018a) whose goal is to predict the precise probability scores for multiple objectives, e.g., click through rate (CTR) and click-through conversion rate (CTCVR). Whereas ME stage is a light-weight model whose goal is to generate a unified ensemble score in a personalized manner. The ensemble score is used to truncate top items to represent for users. Although extensive research has been conducted on MP (MTL) problem in recent years, the ME problem remains more underexplored due to its strong dependence on industrial applications.
Motivation. We ask, what requirements should an ideal ensemble score satisfy? We answer this from two perspectives. On one hand, for each individual objective, to achieve its maximum value (e.g., purchases), the scores must be as consistent as possible with the ranks of its labels (e.g., labels of buy behaviors). On the other hand, for the entire set of multiple objectives, an overall optimum can be achieved only when the scores align with the commonality shared by the majority of objectives. These imply two specific requirements for the design of ME models (see Fig. 1 (b) for conceptual view). R1: Rank consistency. For optimization objective, they should be capable of optimizing towards ranking consistency between ensemble scores and multi-objective labels directly. R2: Commonality modeling. For representation learning, they should be able to model the commonality among multiple objectives. However, we find that existing methods fail to meet these. First, popular methods are optimized by classification or regression tasks (Li et al., 2023; Cao et al., 2025). These tasks are misaligned with the ultimate ranking task (evaluated by AUC) (Tang et al., 2022). Second, although these methods build complex user-specific structure for multi-objective score fusion, none of them consider modeling the underlying commonality across objectives (see Fig. 2 (a)).
@@
Our method. To fill this gap, in this paper, we propose a novel personalized multi-objective ensemble framework HarmonRank, which enables both rank consistency and commonality modelling among multiple objectives (see Fig. 2 (b)). For rank consistency, we firstly reformulate the non-differentiable AUC as a rank sum problem and utilize an advanced differentiable ranking technique to optimize multi-objective AUC in an end-to-end manner. For commonality modeling, we propose to change existing relation-agnostic ensemble architecture to a relation aware one by two steps. First, we align the common parts between multiple objectives and obtain relation-aware representations via self-attention mechanism. Then, we fuse the relation-aware representations of all objectives into the ensemble score in a user-specific manner.
Figure 2. Conceptual comparison between existing methods and our method. (a) Existing methods. User feature conditioned multiple non-linear layers serve as the backbone and multiple binary CE losses serve as the training objectives. (b) Our methods based on two core design insights. (1) Ranking aligned optimization objective towards AUC. (2) Relation-aware architecture with two steps, i.e. commonality modeling, and user-specific ensemble.
@@
  • We propose a novel personalized multi-objective ensemble framework HarmonRank to enable consistency between optimization objective and the ranking task.
  • We propose a relation-aware composite ensemble paradigm to capture the commonality between multiple objectives.
  • We conduct extensive offline experiments on two industrial recommendation datasets. Experimental results demonstrate that our HarmonRank achieves superior performance over all existing state-of-the-art methods. Additionally, HarmonRank contributes 2.6% purchase gain for Kuaishou live-streaming e-commerce platform.
  • We pioneeringly introduce a robustness analysis method for multi-objective ensemble model against label skew. We verify our method outperforms baselines in a large margin on robustness, which makes our method more universal in various industrial scenarios.
@@ 4. HarmonRank
In this section we introduce how HarmonRank fulfill both two aforementioned requirements, rank consistency and commonality modeling. In Sec. 4.1 we elaborate on rank consistency optimization. In Sec. 4.2, we elaborate on design of composite modules for commonality modeling.
4.1. Rank consistency optimization
@@ 4.1.2. Differentiable Ranking
4.2. Commonality modeling
We find prior wisdom consistently overlook the commonality shared among multiple objectives. As shown in Fig. 5, there exist shared parts of varying degrees among objectives. To model the commonalities, we design a relation-aware composite architecture to integrate multiple scores in both relation-agnostic manner and relation-aware manner between objectives. As shown in Fig. 3 (a), our framework is composed of two parallel components: relation-aware module and relation-agnostic module. The output scores of two modules are lately fused into a unified score to compute the AUC loss.
@@ 4.2.1. Relation-aware Module
We align the shared parts between objectives with a relation-aware paradigm by two steps: first align then ensemble. For align step, as shown in Fig. 3(b), we utilize self-attention mechanism (Vaswani et al., 2017) to capture pairwise relations between objectives. Through dynamic computation of relations between objectives, the aligned common parts between them are extracted. Specifically, the softmax normalized dot products between projected objective vectors, $\mathbf{Q}_{r}$ and $\mathbf{K}_{r}$, form the self attention matrix $\mathbf{A}_{r}\in\mathbb{R}^{m\times m}$ between them. This matrix naturally represents the relation strengths between objectives. Then the original score embeddings are integrated into the representations $\mathbf{x}_{r}$ according to the attention matrix.
@@
Although we capture inter-objective commonalities in this way, we still lack clear guidance on how to fuse multiple objectives into a unified score. Fortunately, user-specific personalized information including user profile and context feature, can serve as strong inductive bias to guide this fusion. We cast the personalized information as a query, and search with it over multi-objective scores. We use the searched result as the fused score encodings to obtain the ensemble score. In this way, we implicitly let model answer a question that which objects can best represent the user’s holistic intentions in the current context?. Specifically, we transform personalized information into the query vector $\mathbf{Q}_{p}$ by linear projection, similar to key and value vectors.
@@ 5. Experiments
  • RQ1: How does our model perform on offline datasets?
  • RQ2: How does our module design and hyperparameter choice impact the performance?
  • RQ3: How does our method perform in practical training and inference runtime?
  • RQ4: Can our method bring improvements of A/B test metrics on online product environment?
  • RQ5: How our method perform in severe label skew situation? How our method impact trade-off between objectives? What does the model learn in multi-objective commonality modeling?
@@ 5.1. Implementation Details
Before training ensemble model, we firstly prepare the input multi-objective scores via performing inference with a high-capacity ranking model over both datasets. For Kuaishou-Elive, we use the online in-service model trained incrementally by years of logs. For the public TenRec dataset, we train a standard MMOE (Ma et al., 2018a) ranking model. Based on prepared multi-objective scores, all offline models are trained with 500 epochs with SGD optimizer. In terms of online training model with streaming data, it is usually trained for one epoch for the consideration of run-time cost and one-epoch overfiting risk in ranking stage. This scheme results in low sample efficiency. Since the model of ME stage is much more lightweight than MP stage and no sparse ID feature is used (cause of one-epoch overfitting), we train 20 epochs for each batch of streaming samples, to balance run-time cost and sample efficiency. We set learning rate to 1e-4 and batch size = 10240 for all methods. We train our method and compared methods with the same equal loss weights (all weights set to 1.0) for fair comparison on objective trade-offs.
@@ 5.2. Datasets
Public dataset. TecRec-QKVideo (Yuan et al., 2022) is a popular multi-objective recsys dataset collected from QQ-KAN video feeds platforms. It contains user interaction logs including five type of feedbacks, click, like, share, follow and long view. We sample 2M logs from all dataset, 1M for training, 0.5M for validation and test, respectively. For the need of personalized ensemble, we utilize 2 user profile features, i.e., age, gender, etc.
Private dataset. We build a new dataset called Kuaishou E-Live, which contains both e-commerce feedback (purchase) and user engagement oriented feedbacks (long view, like, follow and comment). We built it by sampling millions of exposure logs at Kuaishou e-commerce platform, which is a pioneering live-streaming e-commerce platform with over 400 million DAUs. We sample 2M logs from one week, 1M for training, 0.5M for validation and test, respectively. We utilize 4 context and user profile features, i.e., the hour of a day, app version, age and gender, to characterize user preference in ensemble learning.
@@ 5.5.2. Model components
5.6. Practical runtime (RQ3)
Training. As shown in Tab. 9, we compare the training speed among different AUC end-to-end losses. The training batch size is set to 10240 for all methods. The ranks of actual training speeds across different methods aligns with their theoretical computational complexity. Due to the pairwise nature of ‘PL’ (Pairwise Logistic), it exhibits the slowest training speed, while AUCM and MBCE methods achieve the fastest training owing to their instance-wise loss computation with $O(n)$ complexity. Although our method’s training speed lies between pairwise and instance-wise approaches, it delivers the best AUC performance across all methods.
Method Practical training speed Theoritical time complexity
MBCE 27.4K/s $O(n)$
PL 24.8K/s $O(n^{2})$
AUCM 26.9K/s $O(n)$
HarmonRank 26.0K/s $O(nlogn)$
Table 5. Comparison of practical training speed.
Inference. We test our practical runtime during serving on an Intel 6230R CPU with NVIDIAT4 GPU. The testing batchsize is 10240. As shown in Tab. 6, our method costs negligible FLOPs and parameters, which is very appropriate for multi-objective ensemble stage.
FLOPs Params GPUMem Latency Throughput
273M 2.4K 2.0GB 4ms 2.5M/s
Table 6. Practical inference cost of our model.
5.7. Online Results (RQ4)
Figure 5. Consistency analysis between the learned attention matrix $\mathbf{r}^{PRED}$ and ground-truth data distribution $r^{GT}$ (label’s Spearman rank correlation.
@@ 5.7. Online Results (RQ4)
Objectives RL Ours
Purchase +1.547% +2.635%
Comment -0.253% +3.034%
Follow +0.302% +0.451%
Watch time -0.238% +0.290%
Like +0.817% +1.673%
@@
5.8. Analysis (RQ5)
5.8.1. Robustness to label skew.
To validate the robustness of our method against label skew, we conduct extensive experiments on the Kuaishou E-live dataset. Among our five objectives, the buy objective exhibits the most severe label skew with a positive-to-negative ratio of $1:10^{3}$. To explore the performance of different methods under harder situations, we artificially strengthen the imbalance of buy objective by down-sampling positive samples to ratios of $1:10^{4}$ and $1:10^{5}$ respectively on training set only, remaining the test set unchanged.
As shown in the Tab. 10, when facing increasingly severe label skew, all methods exhibit expected AUC degradation. Nevertheless, our HarmonRank consistently maintains superior performance across all scenarios. Crucially, the performance drop of our method (-0.3%) is significantly smaller than those of MBCE and PL (9.8% and 4.3%), demonstrating our methods has exceptional robustness to label distribution skew. We attribute this phenomenon to different degrees of dependable on negative samples of these methods. As positive samples become sparser, classifying negative samples becomes increasingly trivial, resulting in diminishing gradients from the overwhelming majority of easy negative examples. Consequently, methods like MBCE and PL that heavily rely on negative samples suffer significant performance degradation. In contrast, HarmonRank inherently rely less on negative samples for its rank-sum formulation, whose loss only computes ranks of positive samples and negative samples are utilized only during the differentiable ranking phase.
[…]
Table 8. Robustness against label skew.
@@ 5.8. Analysis (RQ5)
5.8.3. Inter-objective relation analysis.
To shed light on the factual commonalities among objectives (ground-truth), and which of them can be learned in inter-objective relations modules (predictions), we visualize inter-objective relations in two ways. For the ground-truth relations, we compute Spearman’s rank correlation coefficients $\mathbf{\rho}$ between different objectives over the labels, which measures the monotonic correlations between the ranks of two objective labels.
(22) $r_{mn}^{GT}$ $=\text{Spearmanr}(\mathbf{y}_{m},\mathbf{y}_{n})$
For predicted relations, we visualize the attention weights inferred by our method’s relation module.
(23) $r_{mn}^{PRED}$ $=\mathbf{A}_{r}[m,n]$
As shown in Fig .5, we can observe remarkable similarity between learned attentions $\mathbf{r}^{PRED}$ and empirical label correlations $\mathbf{r}^{GT}$, confirming our method’s capability to accurately capture inter-objective commonalities.